Solve the following Bernoulli's Differential Equations. Show your solutions.



a. dy/dx + (1/3) y = e^x y²



b. x (dy/dx) + y = xy³



c. dy/dx + (2/x) y = -x² cos x y²



d. x²y-x³ (dy/dx) = y² cos x



1
Expert's answer
2021-12-28T12:08:01-0500

a.

"\frac{y'}{y^2}+\frac{1}{3y}=e^x"


"z=1\/y,z'=-y'\/y^2"


"-z'+z\/3=e^x"

"z=uv,z'=u'v+uv'"

"-(u'v+uv')+uv\/3=e^x"

"-u'v+u(v\/3-v')=e^x"


"v\/3-v'=0"

"-u'v=e^x"


"dv\/v=dx\/3"

"lnv=x\/3"

"v=e^{x\/3}"


"-u'e^{x\/3}=e^x"

"u'=-e^{2x\/3}"

"u=-3e^{2x\/3}\/2+c"


"z=e^{x\/3}(-3e^{2x\/3}\/2+c)"


"y=\frac{1}{e^{x/3}(-3e^{2x/3}/2+c)}"


b.

"\frac{xy'}{y^3}+\frac{1}{y^2}=x"


"z=1\/y^2,z'=-2y'\/y^3"


"-xz'\/2+z=x"

"z=uv,z'=u'v+uv'"

"-x(u'v+uv')+2uv=2x"

"-xvu'+u(2v-xv')=2x"


"-vu'=2"

"2v-xv'=0"


"dv\/v=2dx\/x"

"lnv=2lnx"

"v=x^2"


"-x^2u'=2"

"du=-2dx\/x^2"

"u=2\/x+c"

"z=x^2(2\/x+c)"


"y=\frac{1}{x\sqrt{2/x+c}}"


c.

"y'\/y^2+2\/(xy)=-x^2cosx"

"z=1\/y,z'=-y'\/y^2"

"-z'+2z\/x=-x^2cosx"

"z=uv,z'=u'v+uv'"

"-(u'v+uv')+2uv\/x=-x^2cosx"


"-u'v=-x^2cosx"

"2v\/x-v'=0"


"dv\/v=2dx\/x"

"v=x^2"


"u'=cosx"

"u=sinx+c"

"z=x^2(sinx+c)"


"y=\frac{1}{x^2(sinx+c)}"


d.

"x^2\/y-x^3y'\/y^2=cosx"

"z=1\/y,z'=-y'\/y^2"

"zx^2+x^3z'=cosx"

"z=uv,z'=u'v+uv'"

"uvx^2+x^3(u'v+uv')=cosx"


"x^3u'v=cosx"

"vx^2+x^3v'=0"


"dv\/v=-dx\/x"

"lnv=-lnx"

"v=1\/x"


"x^2u'=cosx"


"u=\int cosxdx/x^2=-\frac{i(\Gamma(-1,ix))-\Gamma(-1,-ix)}{2}+c"


"z=\frac{1}{x}(-\frac{i(\Gamma(-1,ix))-\Gamma(-1,-ix)}{2}+c)"


"y=\frac{x}{-\frac{i(\Gamma(-1,ix))-\Gamma(-1,-ix)}{2}+c}"


where "\Gamma(z)" is gamma function:

"\Gamma(z)=\int^{\infin}_0 x^{z-1}e^{-x}dx"